OpenStudy (mallorysipp234):

What is the difference between A=Pe^rt and A=P(1+r/n)^nt?

OpenStudy (peachpi):

A = Pe^(rt) is the formula for continuously compounded interest, where the principle earns interest continuously. A = P(1 + r/n)^(nt) is the formula for periodically compounded interest, where the principle earns interest depending on the period (daily, quarterly, monthly, yearly, etc.)

OpenStudy (mallorysipp234):

so if i have compounded continuously its A=pe^RT and if its monthly for example its the other one? thankyou

OpenStudy (peachpi):

yes. you're welcome